# LLM-Crowdsourced: A Benchmark-Free Paradigm for Mutual Evaluation of Large Language Models

### v1

@@ Introduction
Question Generation Mechanism Design
Mathematics, as a highly formalized language, possesses rigorous reasoning rules and objective standard answers, making it inherently suitable for evaluating the logical reasoning, abstract modeling, and generalization capabilities of LLMs. To mitigate the issue of data contamination caused by static question banks, we have designed a more professional and dynamic mathematical evaluation framework. In the question generation phase (see Figure 3), the question-generating LLM is tasked with acting as a mathematician and formulating professional questions in the mathematical domain. Recognizing that LLMs have encountered substantial mathematical content during training (Achiam et al. 2023), we do not restrict the question type. We require the LLM to identify the most difficult questions based on its prior knowledge, abstract the core principles, and construct an original question accordingly, while also providing a complete reference solution. This mechanism not only increases the novelty and difficulty of the questions but also accentuates the capability differences among LLMs and differentiates between “memorization-based answering” and “reasoning-based answering.”
Evaluation Mechanism Design
Since mathematical questions typically allow for various reasonable solutions, traditional scoring methods struggle to capture different lines of reasoning and are prone to introducing subjective bias from the LLMs. To address this, we employ a relative evaluation mechanism. That is, each LLM evaluates others based on the reference solutions, and the resulting rankings are normalized into final scores. This approach mitigates the subjective bias introduced by absolute scoring standards and promotes more stable and equitable comparisons in multi-LLM evaluations.
@@
Figure 4 shows that the eight mainstream LLMs exhibit significant differences in their performance on mathematical tasks. The figure presents the scores and fluctuation ranges for each LLM in multi-turn question answering, reflecting overall performance and answering stability. It can be observed that LLMs with leading scores, such as gemini-2.5-pro-exp and gpt-4.1, demonstrate balanced performance across various questions, exhibiting strong and stable logical reasoning abilities. In contrast, lower-ranked LLMs, such as o3-mini and o1, show large score fluctuations and evident weaknesses. Overall, the figure intuitively displays the differences in mathematical capabilities among the LLMs and also verifies the effectiveness of this evaluation paradigm in capability differentiation.
@@
Findings 3: LLM evaluation results demonstrate high consistency (robustness). During the evaluation phase, we found that multiple LLMs exhibited high consistency in evaluating the answers of other LLMs. For example, GPT-4.1’s answers were consistently rated by several LLMs as having the clearest logical chain, and the answers from gemini-2.5-pro-exp were frequently deemed superior to the reference answers. Top-k consistency analysis shows that when $k=n/2$, the Top-k value is 74.85%, indicating that most LLMs maintain stable evaluation standards when evaluating other LLMs, thus forming a cross-LLM “consensus judgment standard.” This result also indirectly confirms the robustness and objectivity of our method.
@@
Question Generation Mechanism Design
Programming tasks involve both theoretical complexity and engineering implementation, yet existing benchmarks primarily focus on code output accuracy, seldom addressing question design, which makes it challenging to reflect a LLM’s capability in program synthesis comprehensively. To address this, we propose a professional and dynamic programming evaluation framework. As shown in Figure 5, in each round, the question-generating LLM acts as a senior programming competition question setter, designing professional-level questions within the programming domain. By studying high-level competition questions (such as ACM), the LLM masters structural design and complexity factors of high-quality programming questions. Following this style, the LLM generates a brand-new, non-derivative, and difficult original programming question, and provides the question description, input/output format, constraint specification, reference solution, and complexity analysis. During the answering phase, LLMs can freely choose the programming language, allowing them to leverage their strengths while eliminating biases from artificial constraints.
Evaluation Mechanism Design
To better reflect professionalism and objectivity, all LLMs are required to evaluate the answers of others based on the reference code and evaluation criteria. The evaluation mechanism employs a 100-point scale, which comprehensively considers multiple dimensions, including correctness, execution efficiency, readability, structural clarity, and memory control. The LLMs conduct cross-evaluation, further enhancing the professional interpretability and objectivity of the evaluation results.
@@
Q1: Do the LLMs generate the questions and answers reliable?
We require question-generating LLMs to provide reference answers simultaneously, ensuring each question has a clear, verifiable solution. All questions underwent manual review, with no unsolvable or logically flawed items found. Even if reference answers have flaws, LLMs can independently evaluate answer quality, avoiding reliance on a single reference. With multiple LLMs rotating as questioners, even low-discrimination questions can be diluted through cross-evaluation. The consistent performance of high-scoring LLMs and the Top-k consistency analysis both confirm the stability of the evaluation.
Q2: If the evaluation LLMs may have biases or errors, how can the results be ensured to be objective and reasonable?
We prohibit self-evaluation and adopt a “cross-evaluation + reference answer” mechanism to reduce bias. Even if misjudgments occur, scores are corrected through multiple rounds of cross-evaluation. We perform three independent runs per experiment to mitigate LLM-induced noise, and the scoring results have demonstrated high consistency and robustness. Among the eight selected LLMs, even the weakest outperforms human-crowdsourced on average. Compared with mainstream benchmarks and Chatbot Arena, although there are slight differences in orderings, our results show consistent overall trends, further confirming reliability. Although evaluation LLMs are not perfect, our goal is to explore a scalable, human-free evaluation paradigm.
@@ Conclusion
To overcome the challenges posed by data contamination, black-box operation, and subjective preference present in mainstream evaluation paradigms, such as white-box evaluation, black-box evaluation, and human-crowdsourced evaluation, this paper proposes LLM-Crowdsourced, a novel benchmark-free evaluation paradigm. This approach possesses four core advantages: dynamic, professional, objective, and transparent. Experiments on two classical domains, mathematics and programming, demonstrate that LLM-Crowdsourced can not only effectively distinguish differences in LLM capability, but also identify phenomena that are difficult to detect using traditional methods, such as “memorization-based answering.” The approach exhibits excellent discriminability and generalizability. We have publicly released the complete evaluation framework and experimental code to support community evaluation of new LLMs and to promote the standardization and advancement of LLM evaluation methods.

### v2

@@ Introduction
Question Generation Mechanism Design: Mathematics, as a highly formalized language, possesses rigorous reasoning rules and objective standard answers, making it inherently suitable for evaluating the logical reasoning, abstract modeling, and generalization capabilities of LLMs. To mitigate the issue of data contamination caused by static question banks, we have designed a more professional and dynamic mathematical evaluation framework. In the question generation phase (see Figure 3), the question-generating LLM is tasked with acting as a mathematician and formulating professional questions in the mathematical domain. Recognizing that LLMs have encountered substantial mathematical content during training (Achiam et al. 2023), we do not restrict the question type. We require the LLM to identify the most difficult questions based on its prior knowledge, abstract the core principles, and construct an original question accordingly, while also providing a complete reference solution. This mechanism not only increases the novelty and difficulty of the questions but also accentuates the capability differences among LLMs and differentiates between “memorization-based answering” and “reasoning-based answering.”
Evaluation Mechanism Design: Since mathematical questions typically allow for various reasonable solutions, traditional scoring methods struggle to capture different lines of reasoning and are prone to introducing subjective bias from the LLMs. To address this, we employ a relative evaluation mechanism. Each LLM evaluates others based on the reference solutions, and the rankings are normalized into final scores. This approach mitigates the subjective bias introduced by absolute scoring standards and promotes more stable and equitable comparisons in multi-LLM evaluations.
@@
Figure 4 shows that the eight mainstream LLMs exhibit significant differences in their performance on mathematical tasks. The figure presents the scores and fluctuations for each LLM in multi-turn question answering, reflecting overall performance and answering stability. It can be observed that LLMs with leading scores, such as gemini-2.5-pro-exp and gpt-4.1, demonstrate balanced performance across various questions, exhibiting strong and stable logical reasoning abilities. In contrast, lower-ranked LLMs, such as o3-mini and o1, show large score fluctuations and evident weaknesses. Overall, the figure intuitively displays the differences in mathematical capabilities among the LLMs and verifies the effectiveness of our method in capability differentiation.
@@
Findings 3: LLM evaluation results demonstrate high consistency (robustness). During the evaluation phase, we found that multiple LLMs exhibited high consistency in evaluating the answers of other LLMs. For example, gpt-4.1’s answers were consistently rated by several LLMs as having the clearest logical chain, and the answers from gemini-2.5-pro-exp were frequently deemed superior to the reference answers. Top-k consistency analysis shows that when $k=n/2$, the Top-k value is 74.85%, indicating that most LLMs maintain stable evaluation standards when evaluating other LLMs, thus forming a cross-LLM “consensus judgment standard.” Moreover, the standard deviations of the scores remain low, ranging from 2.04 to 9.87 on a 100-point scale. These results quantitatively demonstrate the robustness and stability of the evaluation results, even with a limited sample size.
@@
Question Generation Mechanism Design: Programming tasks involve both theoretical complexity and engineering implementation, yet existing benchmarks primarily focus on code output accuracy, seldom addressing question design, which makes it challenging to reflect a LLM’s capability in program synthesis comprehensively. To address this, we propose a professional and dynamic programming evaluation framework. As shown in Figure 5, in each round, the question-generating LLM acts as a senior programming competition question setter, designing professional-level questions within the programming domain. By studying high-level competition questions (such as ACM), the LLM masters structural design and complexity factors of high-quality programming questions. Following this style, the LLM generates a brand-new, non-derivative, and difficult original programming question, and provides the question description, input/output format, constraint specification, reference solution, and complexity analysis. During the answering phase, LLMs can freely choose the programming language, allowing them to leverage their strengths while eliminating biases from artificial constraints.
Evaluation Mechanism Design: To better reflect professionalism and objectivity, all LLMs are required to evaluate the answers of others based on the reference code and evaluation criteria. The evaluation mechanism employs a 100-point scale, which comprehensively considers multiple dimensions, including correctness, execution efficiency, readability, structural clarity, and memory control. The LLMs conduct cross-evaluation, further enhancing the professional interpretability and objectivity of the evaluation results.
@@
Q1: Do the LLMs generate the questions and answers reliable? We require question-generating LLMs to provide reference answers simultaneously, ensuring each question has a clear, verifiable solution. All questions and references were manually reviewed by at least three humans, with no unsolvable or logically flawed items found. Even if reference answers have flaws, LLMs can independently evaluate answer quality, avoiding reliance on a single reference. With multiple LLMs rotating as questioners, even low-discrimination questions can be diluted through cross-evaluation. The high Top-k consistency and the low standard deviations both quantitatively confirm the robustness and stability of the evaluation results, even with a limited sample size.
Q2: If the evaluation LLMs may have biases or errors, how can the results be ensured to be objective and reasonable? We prohibit self-evaluation and adopt a “cross-evaluation + reference answer” mechanism to reduce bias. Even if misjudgments occur, scores are corrected through multiple rounds of cross-evaluation. We perform three independent runs per experiment to mitigate LLM-induced noise, and the scoring results have demonstrated high consistency and robustness. We do not claim to offer a perfect solution. Instead, we propose a scalable, human-free evaluation paradigm to explore more objective and reproducible evaluation methods. Despite the exploratory nature and limited sample size, our results already exhibit strong robustness and lay the foundation for future large-scale studies. Among the eight selected LLMs, even the weakest outperforms humans on average. Compared with mainstream benchmarks and Chatbot Arena, although there are slight differences in orderings, our results show consistent overall trends, further confirming reliability.
@@ Conclusion
To overcome the challenges posed by data contamination, black-box operation, and subjective preference present in mainstream evaluation paradigms, such as white-box evaluation, black-box evaluation, and human-crowdsourced evaluation, this paper proposes LLM-Crowdsourced, a novel benchmark-free evaluation paradigm. It integrates four key evaluation criteria: dynamic, transparent, objective, and professional. Experiments on two classical domains, mathematics and programming, demonstrate that LLM-Crowdsourced can not only effectively distinguish differences in LLM capability, but also identify phenomena that are difficult to detect using traditional methods, such as “memorization-based answering.” This method exhibits excellent discriminability and generalizability. We have publicly released the complete evaluation framework and code to support community evaluation of new LLMs and to promote the standardization and advancement of LLM evaluation methods.